OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 2 where f of x equals x plus 3 when x is less than 2 and f of x equals 3 minus x when x is greater than or equal to 2 Would it be the limit Does Not Exist?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

have you plug in 2 into both sides... lets see left side of x=2 we have the function x+3 and the right side of x=2 we have the function 3-x I believe just enter in 2 into both sides 2+3=5 3-2=1 5 is not 1 since 5 is not 1 then the actual limit doesn't exist so you are right

OpenStudy (freckles):

for the limit to have exist 5 would have to be 1

OpenStudy (freckles):

but as you know that is impossible